OpenStudy (calculusxy):

What is: -5/3 x 7/4 =

OpenStudy (calculusxy):

@mathstudent55

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (dannyrod2000):

-35/12

OpenStudy (calculusxy):

I meant 5/3 to be a fraction that is completely negative.

OpenStudy (calculusxy):

Can someone please help me? I am not too good with negatives.

OpenStudy (calculusxy):

@ganeshie8

ganeshie8 (ganeshie8):

-a*b = -(a*b)

ganeshie8 (ganeshie8):

-5/3 x 7/4 = -(5/3 x 7/4)

OpenStudy (calculusxy):

So it will be 35/12 completely negative right?

ganeshie8 (ganeshie8):

-35/12

ganeshie8 (ganeshie8):

(-35)/12 = 35/(-12) = -(35/12)

OpenStudy (calculusxy):

So can you please help with the negatives with addition, subtraction, multiplication, and division.

OpenStudy (calculusxy):

Like can you explain to me them? E.g. 14-(-11) =25?

ganeshie8 (ganeshie8):

-*- = +

ganeshie8 (ganeshie8):

14-(-11) = 14 + 11 = ?

OpenStudy (calculusxy):

And one more question @ganeshie8

ganeshie8 (ganeshie8):

sure ask :)

OpenStudy (calculusxy):

What is the average of 20,6,-5,-8,4,14,-3?

ganeshie8 (ganeshie8):

how many numbers are they ?

OpenStudy (calculusxy):

Can you explain to me the steps?

OpenStudy (calculusxy):

Seven

ganeshie8 (ganeshie8):

add them up, then divide by 7

ganeshie8 (ganeshie8):

average value = (20 + 6 + -5 + -8 + 4 + 14 + -3) / 7

OpenStudy (anonymous):

When in doubt, listen to ganeshie.

OpenStudy (calculusxy):

So let's say it's 26+(-5)= 21 21+ (-8) = 13 13 + 4 = 17 17 + (-18)=-1 -18 + -3 = -21? So I made couple of mistakes in the numbers.

OpenStudy (calculusxy):

Basically if you add two negative numbers will THey also equal negative

ganeshie8 (ganeshie8):

So let's say it's 26+(-5)= 21 21+ (-8) = 13 13 + 4 = 17 `17 + (-18)=-1` -18 + -3 = -21?

ganeshie8 (ganeshie8):

why are you adding -18 in that line ?

OpenStudy (calculusxy):

I made a mistake in the list of the numbers.

OpenStudy (calculusxy):

Anyway, will -18 + -3 equal to -21?

OpenStudy (calculusxy):

I got the answer, thanks for ur help :)

ganeshie8 (ganeshie8):

if you borrow $18 from me today and again borrow $3 from me how much do you owe me ?

ganeshie8 (ganeshie8):

-18 + -3 = -21